2. An exclusive Yoghurt manufacturer sells 4,000 gallons per month at a price of GHS 40 each. When the price is reduced to GHS 30 sales increase to 6,000 gallons per month.

a. Calculate the price elasticity of demand for the Yoghurts over this price range.

b. Is demand elastic, unit elastic or inelastic?

c. Calculate the change in revenue due to the change in price.



1
Expert's answer
2021-05-06T14:24:12-0400

Arc Elasticity, "Arc Ed =[\frac {(Qd2 – Qd1)} {midpoint Qd} ] ÷ [\frac {(P2 – P1)} {midpoint P} ]"


Midpoint "Q_d=\frac {6000+4000} {2} =5,000"


Midpoint Price "=\frac {30+40}{2} =35"


Percentage Change in Quantity demanded ;

"=\frac {6000-4000} {5000} =0.4"


Percentage change in Price ;

"\frac {30-40}{35}=-0.286"


"Arc E_d=\frac {0.4}{-0.286}=1.4"


The demand is elastic .


As the demand is elastic , so fall in price will increase the revenue .